# Learning Constraints from Stochastic Partially-Observed Closed-Loop Demonstrations

### v1

@@ II-B System Level Synthesis (SLS) for Output Feedback
Given a parameter value $\theta\in\mathbb{R}^{d}$, let the set of corresponding safe (resp., unsafe) nominal trajectories and system responses, given below by $\mathcal{S}(\theta)$ (resp., $\mathcal{A}(\theta)$), be defined by:
$\mathcal{S}(\theta)$ $:=\{(\textbf{z},\textbf{v},\Phi):\tilde{\textbf{g}}_{\urcorner{k}}(\textbf{z},\textbf{v},\Phi,\theta)\leq\textbf{0}\},$ (15)
$\mathcal{A}(\theta)$ $:=\mathcal{S}(\theta)^{c}.$ (16)
In particular, $\mathcal{S}(\theta)$ contains the feasible set of (14).
@@
$\tilde{g}_{\beta,k}(\cdot)$ and $\tilde{g}_{\beta,{\urcorner{k}}}(\cdot)$ can be written in closed form for certain constraint functions $g_{\beta,k}(\cdot)$ and $g_{\beta,{\urcorner{k}}}(\cdot)$ and disturbance sets $W$ and $E$ [12, Prop. 4], [13, Prop III.3].
II-C Demonstration Generation Process
We describe how a set of $D$ input-output trajectory demonstrations, denoted $\mathcal{D}:=\{(\textbf{u}^{({d})},\textbf{y}^{({d})})\in\mathbb{R}^{n_{i}T}\times\mathbb{R}^{n_{o}T}:d\in[D]\}$ are generated for constraint learning. First, the demonstrator solves (14) to compute a locally-optimal nominal trajectory $(z^{\star},v^{\star})$ and system response $\Phi^{\star}$, respectively, and computes the corresponding output feedback law $\mathcal{K}^{\star}$ by substituting $\Phi=\Phi^{\star}$ into (12). Then, for each $d\in[D]$, the demonstrator generates the input-output trajectory $(\textbf{u}^{({d})},\textbf{y}^{({d})})\in\mathbb{R}^{n_{i}T}\times\mathbb{R}^{n_{o}T}$, by unrolling the dynamics (1a), and applying output map (1b) and output feedback law (4):
@@ II-C Demonstration Generation Process
$u_{t}^{({d})}$ $=v_{t}^{\star}+\sum_{\tau=0}^{t}K_{t,\tau}^{\star}(y_{\tau}^{({d})}-Cz_{\tau}^{\star}).$ (17c)
Above, $(\textbf{w}^{({d})},\textbf{e}^{({d})})\in W^{T+1}\times E^{T+1}$ describes the dynamics and output noise realizations for the $d$-th input-output trajectory, for each $d\in[D]$. Finally, the constraint learner receives the perturbed demonstration set $\tilde{D}:=\{(\tilde{\textbf{u}}^{({d})},\tilde{\textbf{y}}^{({d})})\in\mathbb{R}^{n_{i}T}\times\mathbb{R}^{n_{o}T}:d\in[D]\}$, with:
@@
where $(\bm{\delta}_{u}^{({d})},\bm{\delta}_{y}^{({d})})\in\mathbb{R}^{n_{i}T}\times\mathbb{R}^{n_{o}T}$ represent possible transmission errors that may corrupt the constraint learner’s reception of the generated input-output demonstrations.
@@
Above, we considered multiple input-output trajectories generated from different rollouts of the same nominal trajectory using the same output feedback law. However, our methods also extend to learn constraints from multiple nominal trajectories and feedback laws.
II-D Problem Statement
We aim to infer the unknown constraint parameter $\theta^{\star}$ from the perturbed input-output demonstration set $\tilde{D}$, as described above, as well as knowledge of the system equations (1), costs $J$ and $\tilde{J}$, and constraints $g_{\beta,k}$, $g_{\beta,{\urcorner{k}}}$, $\tilde{g}_{\beta,k}$, and $\tilde{g}_{\beta,{\urcorner{k}}}$. We also aim to learn nominal trajectories and output feedback laws that are guaranteed to lie within the safe set $\mathcal{S}(\theta)$ despite worst-case dynamics and output noise realizations.
III Methods
We present our method for learning the nominal trajectory, output feedback law, and unknown constraints (Sec. III-A) and prove that under zero transmission noise and mild conditions, the recovered values are accurate (Sec. III-B). We also perform sensitivity analysis to prove that learning error grows linearly with transmission noise magnitude (Sec. III-C).
III-A Recovering the Nominal Trajectory, Output Feedback, and Constraint Parameters
@@
We define $\tilde{\textbf{U}}\in\mathbb{R}^{n_{i}T\times(D-1)}$, $\tilde{\textbf{Y}}\in\mathbb{R}^{n_{o}T\times(D-1)}$ and compute $\tilde{\mathcal{K}}$ as:
@@
$\tilde{\mathcal{K}}:=\tilde{\textbf{U}}\tilde{\textbf{Y}}^{\dagger}$. (23)
We then use (11) to compute the system response $\tilde{\Phi}$ corresponding to $\tilde{\mathcal{K}}$.
@@
We then estimate $(\textbf{z}^{\star},\textbf{v}^{\star})$ from $\tilde{\mathcal{K}}$ as follows [^2] For the case of nonlinear dynamics, see Remark 6.:
@@
Above, we average over the terms $\{\tilde{\textbf{u}}^{({d})}-\tilde{\mathcal{K}}\tilde{\textbf{y}}^{({d})}:d\in[D]\}$ to reduce the impact of the transmission errors $\{\bm{\delta}_{u}^{({d})},\bm{\delta}_{y}^{({d})}:d\in[D]\}$ on the recovery of $(\textbf{z}^{\star},\textbf{v}^{\star})$.
Finally, we infer the unknown constraint parameter $\theta^{\star}$. For notational ease, given $\textbf{z},\textbf{v},\Phi$, we define $\bm{\eta}:=(\textbf{z},\textbf{v},\Phi)$. We also define $\bm{\nu}_{k}$, $\bm{\lambda}_{k}$, and $\bm{\lambda}_{\urcorner{k}}$ to denote the Lagrange multipliers associated with (14b), (14c), and (14d), respectively. The KKT conditions of (14) are then given by:
$\textbf{h}(\bm{\eta})=0,\hskip 14.22636pt\textbf{g}_{k}(\bm{\eta})\leq 0,\hskip 14.22636pt\textbf{g}_{\urcorner{k}}(\bm{\eta},\theta)\leq 0,$ (25a)
$\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}}\geq 0,$ (25b)
$\bm{\lambda}_{k}\odot\textbf{g}_{k}(\bm{\eta})=0,\hskip 14.22636pt\bm{\lambda}_{\urcorner{k}}\odot\textbf{g}_{\urcorner{k}}(\bm{\eta},\theta)=0,$ (25c)
@@
$\hskip 5.69054pt+\bm{\lambda}_{\urcorner{k}}^{\top}\nabla_{\bm{\eta}}\textbf{g}_{\urcorner{k}}(\bm{\eta},\theta)+\bm{\nu}^{\top}\nabla_{\bm{\eta}}\textbf{h}(\bm{\eta})=0.$
We denote by $\text{KKT}(\bm{\eta})$ the set of all $(\theta,\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}},\bm{\nu})$ satisfying (25). Since the true nominal trajectory and output response computed by the demonstrator, which we denote by $\bm{\eta}^{\star}:=(\textbf{z}^{\star},\textbf{v}^{\star},\Phi^{\star})$, minimizes (14), there exist Lagrange multiplier values $\bm{\lambda}_{k}^{\star},\bm{\lambda}_{\urcorner{k}}^{\star},\bm{\nu}^{\star}$ such that $(\theta^{\star},\bm{\lambda}_{k}^{\star},\bm{\lambda}_{\urcorner{k}}^{\star},\bm{\nu}^{\star})$ solves the following feasibility problem with $\bm{\eta}=\bm{\eta}^{\star}$:
@@
Conversely, solutions to (26) with $\bm{\eta}=\bm{\eta}^{\star}$ describe parameter values $\theta$ consistent with the optimality of $\bm{\eta}^{\star}$ with respect to (14), and thus inform the true parameter value $\theta^{\star}$. For convenience, given any $\bm{\eta}=(\textbf{z},\textbf{v},\Phi)$, we denote by $\mathcal{F}(\bm{\eta})\subseteq\mathbb{R}^{d}$ the set of KKT-compatible parameter values, i.e.,
@@
III-B Learning Guarantees Under Zero Transmission Error
We now establish sufficient conditions under which $\bm{\eta}^{\star}$, the ground truth nominal trajectories and system response, can be accurately recovered.
@@ III-B Learning Guarantees Under Zero Transmission Error
$\tilde{\textbf{Y}}$ is full row rank and $\Gamma$ (28) is invertible:
$\Gamma:=\begin{bmatrix}I-\mathcal{Z}\mathcal{A}&-\mathcal{Z}\mathcal{B}\\ -\mathcal{K}^{\star}\mathcal{C}&I\end{bmatrix}\in\mathbb{R}^{(n+n_{i})T\times(n+n_{i})T}.$ (28)
In words, Assumption 1 states that the provided demonstrations $\tilde{\textbf{Y}}$ are sufficiently rich for recovering $\mathcal{K}^{\star}$, and that $\mathcal{K}^{\star}$ satisfies a mild regularity condition. In particular, if $\bm{\delta}_{u}^{({d})}$ and $\bm{\delta}_{y}^{({d})}$ are i.i.d. across indices $d\in[D]$, then $\tilde{\textbf{Y}}$ has full row rank with zero mean whenever $D\geq n_{o}T+1$. Thm. 1 below shows that in the absence of transmission noise, Assumption 1 is sufficient to guarantee the accurate recovery of $\bm{\eta}^{\star}$.
Theorem 1
@@
which can be stacked across all $d\in[2,D]$ as $\tilde{\textbf{U}}=\mathcal{K}^{\star}\tilde{\textbf{Y}}$. Since $\tilde{\textbf{Y}}$ has full row rank, $\tilde{\textbf{Y}}\tilde{\textbf{Y}}^{\dagger}=I_{n_{o}T}$, so $\mathcal{K}^{\star}=\tilde{\textbf{U}}\tilde{\textbf{Y}}^{\dagger}$. Together with (23) and (11), this confirms that $\mathcal{K}^{\star}=\tilde{\mathcal{K}}$ and $\Phi^{\star}=\tilde{\Phi}$. Then, from (24) and preceding equations, we have under Assumption 1 that $(\tilde{\textbf{z}},\tilde{\textbf{v}})=(\textbf{z}^{\star},\textbf{v}^{\star})$. ∎
Remark 6
If the dynamics (1a) are nonlinear, we can still recover the output feedback law via (23) (since the output map is assumed LTV). However, we would apply nonlinear regression techniques to fit the input-output demonstrations, while enforcing the dynamics $z_{t+1}^{\star}=f_{t}(z_{t}^{\star},v_{t}^{\star})$, instead of applying the procedure described in Sec. III-A. Although the nominal trajectories would no longer be guaranteed to be accurately recovered (in the sense of Thm. 1), we find that in simulation, we still recover the true nominal trajectories with minimal error when enough demonstrations are provided.
@@
$\mathcal{G}_{s}(\bm{\eta}^{\star})$ $:=\bigcap_{\theta\in\mathcal{F}(\bm{\eta}^{\star})}\big\{\bm{\eta}^{\prime}:\textbf{g}(\bm{\eta}^{\prime},\theta)\leq 0\big\}=\bigcap_{\theta\in\mathcal{F}(\bm{\eta}^{\star})}\mathcal{S}(\theta),$
@@
$\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})$ $:=\bigcap_{\theta\in\mathcal{F}(\bm{\eta}^{\star})}\big\{\bm{\eta}^{\prime}:\textbf{g}(\bm{\eta}^{\prime},\theta)\leq 0\big\}^{c}=\bigcap_{\theta\in\mathcal{F}(\bm{\eta}^{\star})}\mathcal{A}(\theta).$
@@
Below, Thm. 2 states that $\mathcal{G}_{s}(\bm{\eta}^{\star})$ is guaranteed to inner-approximate (i.e., conservatively estimate) $\mathcal{S}(\theta^{\star})$, the set of nominal trajectories and system responses $\bm{\eta}$ that are safe with respect to the ground truth inequality constraints. Similarly, $\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})$ always inner-approximates $\mathcal{A}(\theta^{\star})$.
Theorem 2
@@
$\mathcal{G}_{s}(\bm{\eta}^{\star})\subseteq\mathcal{S}(\theta^{\star})$ and $\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})\subseteq\mathcal{A}(\theta^{\star})$.
Proof:
@@
By the definitions of $\mathcal{G}_{s}(\bm{\eta}^{\star})$ and $\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})$, it suffices to prove $\theta^{\star}\in\mathcal{F}(\bm{\eta}^{\star})$. Since $\bm{\eta}^{\star}$ solves (14), there exist Lagrange multipliers $\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}},\bm{\nu}$ such that $(\theta^{\star},\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}},\bm{\nu})\in\text{KKT}(\bm{\eta}^{\star})$. Thus, by definition of $\mathcal{F}(\bm{\eta}^{\star})$, we have $\theta^{\star}\in\mathcal{F}(\bm{\eta}^{\star})$, so $\mathcal{G}_{s}(\bm{\eta}^{\star})\subseteq\mathcal{S}(\theta^{\star})$ and $\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})\subseteq\mathcal{A}(\theta^{\star})$. ∎
Remark 7
With respect to many real-world constraints (e.g., collision-avoidance), (26) can be reformulated as a mixed-integer linear program (MILP) or a mixed-integer bilinear program (MIBLP) [1, Sec. 4B, 4D], and solved via off-the-shelf solvers (e.g., Gurobi) [17].
III-C Sensitivity Analysis With Respect to Transmission Error

### v2

@@ 2.2 System Level Synthesis (SLS) for Output Feedback
Given a parameter value $\theta\in\mathbb{R}^{d}$, let the set of corresponding safe (resp., unsafe) nominal trajectories and system responses, given below by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{S}}(\theta)$ (resp., ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{A}}(\theta)$), be defined by:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{S}}(\theta)$ $:=\{(\textbf{z},\textbf{v},\Phi):\tilde{\textbf{g}}_{\urcorner{k}}(\textbf{z},\textbf{v},\Phi,\theta)\leq\textbf{0}\},$ (15)
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{A}}(\theta)$ $:={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{S}}(\theta)^{c},$ (16)
where $(\cdot)^{c}$ indicates the complement of a set. In particular, ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{S}}(\theta)$ contains the feasible set of (14).
@@
$\tilde{g}_{\beta,k}(\cdot)$ and $\tilde{g}_{\beta,{\urcorner{k}}}(\cdot)$ can be written in closed form for certain constraint functions $g_{\beta,k}(\cdot)$ and $g_{\beta,{\urcorner{k}}}(\cdot)$ and disturbance sets $W$ and $E$ [8, Prop. 4], [9, Prop III.3].
2.3 Demonstration Generation Process
We describe how a set of $D$ input-output trajectory demonstrations $\mathcal{D}:=\{(\textbf{u}^{({d})},\textbf{y}^{({d})})\in{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{R}^{(n_{i}+n_{o})(T+1)}}:d\in[D]\}$ are generated for constraint learning. First, the demonstrator solves (14) to compute a locally-optimal nominal trajectory $(z^{\star},v^{\star})$ and system response $\Phi^{\star}$, respectively, and computes the corresponding output feedback law $\mathcal{K}^{\star}$ by substituting $\Phi=\Phi^{\star}$ into (12). Then, $\forall d\in[D]$, the demonstrator generates $(\textbf{u}^{({d})},\textbf{y}^{({d})})$ by unrolling the dynamics (1a), and applying output map (1b) and output feedback law (4):
@@ 2.3 Demonstration Generation Process
$u_{t}^{({d})}$ $=v_{t}^{\star}+\textstyle\sum_{\tau=0}^{t}K_{t,\tau}^{\star}(y_{\tau}^{({d})}-Cz_{\tau}^{\star}).$ (17c)
Above, $(\textbf{w}^{({d})},\textbf{e}^{({d})})\in W^{T+1}\times E^{T+1}$ describes the dynamics and output noise realizations for the $d$-th input-output trajectory, for each $d\in[D]$. Finally, the constraint learner receives the perturbed demonstration set $\tilde{D}:=\{(\tilde{\textbf{u}}^{({d})},\tilde{\textbf{y}}^{({d})})\in\mathbb{R}^{n_{i}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}}\times\mathbb{R}^{n_{o}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}}:d\in[D]\}$, with:
@@
where $(\bm{\delta}_{u}^{({d})},\bm{\delta}_{y}^{({d})})\in\mathbb{R}^{n_{i}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}}\times\mathbb{R}^{n_{o}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}}$ represent possible transmission errors that may corrupt the constraint learner’s reception of the generated input-output demonstrations.
@@
Our methods also extend to learn constraints from input-output trajectories generated using multiple nominal trajectories and feedback laws, provided the learner is aware which input-output trajectories are identified with which nominal trajectory and feedback law.
2.4 Problem Statement
We aim to infer the unknown constraint parameter $\theta^{\star}$ from the perturbed demonstration set $\tilde{D}$ described above, system equations (1), cost $J$, and constraints $g_{\beta,k}$, $g_{\beta,{\urcorner{k}}}$, $\tilde{g}_{\beta,k}$, and $\tilde{g}_{\beta,{\urcorner{k}}}$. We also aim to learn nominal trajectories and output feedback laws that are guaranteed to lie within the safe set ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{S}}(\theta)$ despite worst-case dynamics and output noise outcomes.
3 Methods
3.1 Recovering the Nominal Trajectory, Output Feedback, and Constraint Parameters
@@
We define $\tilde{\textbf{U}}\in\mathbb{R}^{n_{i}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}\times(D-1)}$, $\tilde{\textbf{Y}}\in\mathbb{R}^{n_{o}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}\times(D-1)}$ and compute $\tilde{\mathcal{K}}$ as:
@@
$\tilde{\mathcal{K}}$ $:=\tilde{\textbf{U}}\tilde{\textbf{Y}}^{\dagger}$, (23)
where $\dagger$ denotes the Moore-Penrose pseudoinverse. We then compute the associated system response $\tilde{\Phi}$ via (11).
@@
We then estimate $(\textbf{z}^{\star},\textbf{v}^{\star})$ from $\tilde{\mathcal{K}}$ as follows [^2] For the case of nonlinear dynamics, see Remark 3.4.:
@@
Above, we average over the terms $\{\tilde{\textbf{u}}^{({d})}-\tilde{\mathcal{K}}\tilde{\textbf{y}}^{({d})}:d\in[D]\}$ to reduce the impact of the transmission errors $\{\bm{\delta}_{u}^{({d})},\bm{\delta}_{y}^{({d})}:d\in[D]\}$ on the recovery of $(\textbf{z}^{\star},\textbf{v}^{\star})$. Finally, we infer the unknown constraint parameter $\theta^{\star}$. To ease notation, given $\textbf{z},\textbf{v},\Phi$, we define $\bm{\eta}:=(\textbf{z},\textbf{v},\Phi)$. We also define $\bm{\nu}$, $\bm{\lambda}_{k}$, and $\bm{\lambda}_{\urcorner{k}}$ to denote the Lagrange multipliers for (14b), (14c), and (14d), respectively. The KKT conditions of (14) are then:
$\textbf{h}(\bm{\eta})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\textbf{0}},\hskip 14.22636pt\textbf{g}_{k}(\bm{\eta})\leq\textbf{0},\hskip 14.22636pt\textbf{g}_{\urcorner{k}}(\bm{\eta},\theta)\leq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\textbf{0}},$ (25a)
$\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}}\geq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\textbf{0}},$ (25b)
$\bm{\lambda}_{k}\odot\textbf{g}_{k}(\bm{\eta})=\textbf{0},\hskip 14.22636pt\bm{\lambda}_{\urcorner{k}}\odot\textbf{g}_{\urcorner{k}}(\bm{\eta},\theta)={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\textbf{0}},$ (25c)
@@
$\hskip 5.69054pt+\bm{\lambda}_{\urcorner{k}}^{\top}\nabla_{\bm{\eta}}\textbf{g}_{\urcorner{k}}(\bm{\eta},\theta)+\bm{\nu}^{\top}\nabla_{\bm{\eta}}\textbf{h}(\bm{\eta})={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\textbf{0}^{\top}},$
where $\odot$ denotes element-wise multiplication. We denote by $\text{KKT}(\bm{\eta})$ the set of all $(\theta,\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}},\bm{\nu})$ satisfying (25). Since the true nominal trajectory and output response computed by the demonstrator, which we denote by $\bm{\eta}^{\star}:=(\textbf{z}^{\star},\textbf{v}^{\star},\Phi^{\star})$, minimizes (14), there exist Lagrange multiplier values $\bm{\lambda}_{k}^{\star},\bm{\lambda}_{\urcorner{k}}^{\star},\bm{\nu}^{\star}$ such that $(\theta^{\star},\bm{\lambda}_{k}^{\star},\bm{\lambda}_{\urcorner{k}}^{\star},\bm{\nu}^{\star})$ solves the following feasibility problem with $\bm{\eta}=\bm{\eta}^{\star}$:
@@
Conversely, solutions to (26) with $\bm{\eta}=\bm{\eta}^{\star}$ describe parameter values consistent with the optimality of $\bm{\eta}^{\star}$ with respect to (14); the ground truth $\theta^{\star}$ is one such parameter value. Concretely, given any $\bm{\eta}=(\textbf{z},\textbf{v},\Phi)$, we denote by $\mathcal{F}(\bm{\eta})\subseteq\mathbb{R}^{d}$ the set of KKT-compatible parameter values, i.e.,
@@
3.2 Learning Guarantees Under Zero Transmission Error
First, we prove that the following matrix $\Gamma$ is invertible:
$\Gamma:=\begin{bmatrix}I-\mathcal{Z}\mathcal{A}&-\mathcal{Z}\mathcal{B}\\ -\mathcal{K}^{\star}\mathcal{C}&I\end{bmatrix}$ (28)
Lemma 1
When $f_{t}(\cdot)$ is LTV, $\Gamma$ is invertible.
Proof 3.1.
By the Schur complement, $\det(\Gamma)=\det(I-\mathcal{Z}\mathcal{A})\det(I-\mathcal{Z}(\mathcal{A}+\mathcal{B}\mathcal{K}^{\star}\mathcal{C}))$. Since $\mathcal{A}$, $\mathcal{B}$, $\mathcal{K}^{\star}$, $\mathcal{C}$ are lower block triangular and $\mathcal{Z}$ is the downshift operator, $\mathcal{Z}\mathcal{A}$ and $\mathcal{Z}(\mathcal{A}+\mathcal{B}\mathcal{K}^{\star}\mathcal{C})$ are both strictly lower block triangular. Thus, $\det(\Gamma)=1$, so $\Gamma$ is invertible.
We present a condition under which $\bm{\eta}^{\star}$, the true nominal trajectories and system response, can be accurately recovered.
@@ 3.2 Learning Guarantees Under Zero Transmission Error
$\tilde{\textbf{Y}}$ has full row rank.
In words, Assumption 1 states that the demonstrations $\tilde{\textbf{Y}}$ are sufficiently rich for recovering $\mathcal{K}^{\star}$, and that $\mathcal{K}^{\star}$ satisfies a mild regularity condition. In particular, if $\{\bm{\delta}_{u}^{({d})}:d\in[D]\}$, $\{\bm{\delta}_{y}^{({d})}:d\in[D]\}$ are i.i.d. and independent of $\{w_{t}^{({d})},e_{t}^{({d})}:t\in[T],d\in[D]\}$, then $\tilde{\textbf{Y}}$ has full row rank whenever $D\geq n_{o}(T+1)+1$. Thm. 3.2 below shows that under zero transmission noise, Assumption 1 is sufficient to guarantee the accurate recovery of $\bm{\eta}^{\star}$.
Theorem 3.2.
@@
which can be stacked across all $d\in[2,D]$ as $\tilde{\textbf{U}}=\mathcal{K}^{\star}\tilde{\textbf{Y}}$. Since $\tilde{\textbf{Y}}$ has full row rank, $\tilde{\textbf{Y}}\tilde{\textbf{Y}}^{\dagger}=I_{n_{o}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}(T+1)}}$, so $\mathcal{K}^{\star}=\tilde{\textbf{U}}\tilde{\textbf{Y}}^{\dagger}$. Together with (23) and (11), this confirms that $\mathcal{K}^{\star}=\tilde{\mathcal{K}}$ and $\Phi^{\star}=\tilde{\Phi}$. Then, from (24) and preceding equations, we have under Assumption 1 that $(\tilde{\textbf{z}},\tilde{\textbf{v}})=(\textbf{z}^{\star},\textbf{v}^{\star})$.
Remark 3.4.
If the dynamics (1a) are nonlinear, we can still recover the output feedback law via (23) (since the output map is assumed LTV). However, we would apply nonlinear regression techniques to fit the input-output demonstrations, while enforcing the dynamics $z_{t+1}^{\star}=f_{t}(z_{t}^{\star},v_{t}^{\star})$, instead of applying the procedure described in Sec. 3.1. Although the nominal trajectories would no longer be guaranteed to be accurately recovered (in the sense of Thm. 3.2), we find that in simulation, we still recover the true nominal trajectories with minimal error when enough demonstrations are provided.
@@
$\mathcal{G}_{s}(\bm{\eta}^{\star})$ $:=\textstyle\bigcap_{\theta\in\mathcal{F}(\bm{\eta}^{\star})}\big\{\bm{\eta}^{\prime}:\textbf{g}(\bm{\eta}^{\prime},\theta)\leq 0\big\}=\bigcap_{\theta\in\mathcal{F}(\bm{\eta}^{\star})}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{S}}(\theta),$
@@
$\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})$ $:=\textstyle\bigcap_{\theta\in\mathcal{F}(\bm{\eta}^{\star})}\big\{\bm{\eta}^{\prime}:\textbf{g}(\bm{\eta}^{\prime},\theta)\leq 0\big\}^{c}=\bigcap_{\theta\in\mathcal{F}(\bm{\eta}^{\star})}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{A}}(\theta).$
@@
Below, Thm. 3.5 states that $\mathcal{G}_{s}(\bm{\eta}^{\star})$ is guaranteed to inner-approximate (i.e., conservatively estimate) ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{S}}(\theta^{\star})$, the set of nominal trajectories and system responses $\bm{\eta}$ that are safe with respect to the ground truth inequality constraints. Similarly, $\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})$ always inner-approximates ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{A}}(\theta^{\star})$.
Theorem 3.5.
@@
$\mathcal{G}_{s}(\bm{\eta}^{\star})\subseteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{S}}(\theta^{\star})$ and $\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})\subseteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{A}}(\theta^{\star})$.
Proof 3.6.
@@
By the definitions of $\mathcal{G}_{s}(\bm{\eta}^{\star})$ and $\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})$, it suffices to prove $\theta^{\star}\in\mathcal{F}(\bm{\eta}^{\star})$. Since $\bm{\eta}^{\star}$ solves (14), there exist Lagrange multipliers $\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}},\bm{\nu}$ such that $(\theta^{\star},\bm{\lambda}_{k},\bm{\lambda}_{\urcorner{k}},\bm{\nu})\in\text{KKT}(\bm{\eta}^{\star})$. Thus, by definition of $\mathcal{F}(\bm{\eta}^{\star})$, we have $\theta^{\star}\in\mathcal{F}(\bm{\eta}^{\star})$, so $\mathcal{G}_{s}(\bm{\eta}^{\star})\subseteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{S}}(\theta^{\star})$ and $\mathcal{G}_{\urcorner{s}}(\bm{\eta}^{\star})\subseteq{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathbb{A}}(\theta^{\star})$.
3.3 Sensitivity Analysis With Respect to Transmission Error